# The Gauge Picture of Quantum Dynamics

### 03 Jan 2024 v4

@@ Abstract
I Introduction
Locality is of fundamental importance in theoretical physics. The observable physics of quantum dynamics is local if the Hamiltonian is geometrically local [^1] Geometric locality is not to be confused with the weaker notion of $k$-locality, for which an operator is $k$-local if it acts on at most $k$ qubits anywhere in space., i.e. if the Hamiltonian is a sum of local operators. An operator (other than the Hamiltonian) is said to be local if it only acts on a small region of space. But locality is not explicit in the Schrödinger picture because the wavefunction is global in the sense that it can not be associated with any local region of space, and the time dynamics of the wavefunction globally depends on all Hamiltonian terms. On the other hand, locality is explicit in the Heisenberg picture, for which the time dynamics of local operators only depends on nearby local operators (when the Hamiltonian is local). Deutsch and Hayden (2000) This motivates us to ask: Is it possible to modify the Schrödinger picture to make locality explicit in the equations of motion?
Gauge theory is another fundamental concept in theoretical physics. Gauge theory is the foundation of the Standard Model of particle physics and is also used to describe exotic phases of condensed matter Levin and Wen (2005); Senthil et al. (2004). An important tool that gauge theory provides is the gauging process, in which one promotes a global symmetry into a local gauge symmetry by coupling the original model to gauge fields. For example, a scalar field theory $\mathcal{L}=\frac{1}{2}(\partial_{\mu}\phi)^{2}$ is invariant under a global $U(1)$ symmetry $\phi(x)\to\phi(x)+\lambda$. By coupling $\phi$ to a gauge field $A_{\mu}$ as in $\mathcal{L}_{\text{gauged}}=\frac{1}{2}(\partial_{\mu}\phi-A_{\mu})^{2}$, the global symmetry is promoted to a local gauge symmetry where $\phi(x)\to\phi(x)+\lambda(x)$ and $A_{\mu}(x)\to A_{\mu}(x)+\partial_{\mu}\lambda(x)$. Gauging more exotic symmetries leads to more exotic physics; e.g. gauging spatial symmetries can lead to gravity and gauging fractal symmetries can result in fracton topological order Yoshida (2013). In quantum mechanics, expectation values $\braket{\psi}{A}{\psi}$ are invariant under a global unitary transformation acting on the wavefunction |$\ket{\psi}\to U\ket{\psi}$ and operators $A\to UAU^{\dagger}$. Although this transformation is typically viewed as a global invariance rather than a global symmetry, we can still ask: Is it possible to gauge the global unitary invariance in quantum mechanics? And what are the consequences of doing so?
@@ I Introduction
The Schrödinger and Heisenberg pictures are related by a time-dependant unitary transformation that acts on the wavefunctions and operators. Our new picture of quantum dynamics is related to the Schrödinger and Heisenberg pictures via a local gauge transformation. We describe these derivations in detail in Sec. II and Sec. III. The new local equations of motion are given in Eq. (19), and the local gauge transformation is given in Eq. (22).
In Sec. IV, we show that the gauge picture local wavefunction associated with a patch is equivalent to the interaction picture wavefunction when the interaction Hamiltonian is the sum over Hamiltonian terms that have some support on that patch. To gain intuition, in Sec. V we consider the example of a quantum circuit in the gauge picture. In Sec. VI, we describe the measurement process in the gauge picture. Although local unitary dynamics are explicitly local in the gauge picture, the affect of local measurement is not explicitly local in the gauge picture (similar to the Schrödinger picture). In Sec. VII, we generalize spatial locality in the gauge picture to e.g. locality in local particle number or local energy density.
II Schrödinger to Heisenberg Picture
@@ III Gauge Picture
III.1 Solving the Equations of Motion
@@ III.2 Generalized Hamiltonians
IV Local Interaction Picture
@@
V Quantum Circuits
(a) Schrödinger: |$\ket{\psi^{\text{S}}(t_{\text{f}})}=U\ket{\psi_{0}}$
(b) Heisenberg: $A^{\text{H}}(t_{\text{f}})=UA(0)\,U^{\dagger}$
(c) local Heisenberg: $A_{I}^{\text{H}}(t_{\text{f}})=UA_{I}(0)U^{\dagger}$
@@ VIII Conclusion
VIII.1 New Approximation Technique
Simulating the gauge picture dynamics on a computer can be achieved using a computational time that scales as $O(n_{\text{U}}N^{3}T_{\text{f}}/\delta_{\text{t}})$, where $n_{\text{U}}$ is the number of connections (which typically scales with the system size), $N$ is the Hilbert space dimension, $N^{3}$ is the time complexity for $N\times N$ matrix multiplication in practice [^2] For extremely large $N$, an $N^{\omega}$ time complexity for matrix multiplication is in principle possible for some $32\leq\omega<$ 3 Hartnett (2021); Strassen (1969)., $T_{\text{f}}$ is the final time, and $\delta_{\text{t}}$ is the time step used for numerical integration. [^3] The time complexity is dominated by the time it takes to multiply a connection $U_{IJ}$ with another $N\times N$ matrix in Eqs. (19) and (20), which must be done $O(n_{\text{U}}T_{\text{f}}/\delta_{\text{t}})$ times. Although this is more expensive than simulating the Schrödinger picture, in another work Slagle (2023) we show that the gauge picture facilitates a new kind of approximation scheme for improved computational efficiency.
@@ VIII.1 New Approximation Technique
Figure 4: (a) A matrix product state in canonical form, where the yellow tensors are isometry tensors. The red tensor can be viewed as a local wavefunction in a truncated Hilbert space. (b) Removing the red tensor from the matrix product state yields the operator that maps the full Hilbert space (bottom) to the truncated Hilbert space (top). For more details, see Appendix B of Ref. Slagle (2023) (from which this figure was adapted).
@@
The gauge picture allows a more efficient Hilbert space truncation where a different subspace of states is kept for different patches of space. The utility of a position-dependent Hilbert space truncation is exemplified by matrix product states Orús (2014) (and isometric tensor networks Zaletel and Pollmann (2020)), which are efficient tensor network representations of wavefunctions with area-law entanglement Orús (2014). Figure 4 (adapted from Ref. Slagle (2023)) shows how a matrix product state can be viewed as a wavefunction in a position-dependent truncated Hilbert space. Although position-dependent truncated Hilbert spaces are already used by matrix product states and isometric tensor networks, the gauge picture provides a new way to approximately time evolve wavefunctions using position-dependent truncated Hilbert spaces. Compared to tensor network methods, the gauge picture provides significantly simpler quantum dynamics algorithms for models in two or more spatial dimensions and for models with fermions. See Ref. Slagle (2023) for a more thorough exploration of this approach.
VIII.2 New Deformation of Quantum Mechanics
@@
The gauge picture agrees with the Schrödinger picture when there is no curvature in the gauge connections, i.e. when Eq. (14) is satisfied. Adding curvature to the initial conditions [i.e. adding violations of Eq. (14)] yields a new kind of deformation of quantum theory, which could yield testable experimental predictions for new physics beyond quantum theory. Considering gauge curvature deformations of quantum theory may be appealing since gauge curvature has been a major theme in the Standard Model of particle physics. Furthermore, this deformation explicitly maintains locality, in contrast to Weinberg’s nonlinear deformations of quantum theory Weinberg (1989) which allows supraluminal communications Gisin (1990); Polchinski (1991). However, future research is needed to determine if testable predictions Slagle (2021) can be made that are consistent with previous experiments.
VIII.3 Other Future Directions
@@
Other intriguing future directions include: (1) It is not clear how to describe quantum channels or the Lindblad master equation in the gauge picture. The obstruction is that we used a unitary operator ($\widetilde{U}_{I}$) to cancel out distant unitary dynamics in order to achieve locality; but unitary operators can not cancel out the non-unitary dynamics of quantum channels or the Lindbladian. (2) Does the gauge picture have useful applications for understanding exotic dynamical many-body phenomena, such as out-of-time-ordered correlation functions Swingle (2018); García-Mata et al. (2022), many body localization Nandkishore and Huse (2015); Abanin et al. (2019), or the speed of information propagation Nachtergaele and Sims (2011); Roberts and Swingle (2016); Wang and Hazzard (2020)?
Acknowledgements.

### 15 Mar 2024 v5

@@ Abstract
1 Introduction
Locality is of fundamental importance in theoretical physics. The observable physics of quantum dynamics is local if the Hamiltonian is geometrically local [^1] Geometric locality is not to be confused with the weaker notion of $k$-locality, for which an operator is $k$-local if it acts on at most $k$ qubits anywhere in space., i.e. if the Hamiltonian is a sum of local operators. An operator (other than the Hamiltonian) is said to be local if it only acts on a small region of space. But locality is not explicit in the Schrödinger picture because the wavefunction is global in the sense that it can not be associated with any local region of space, and the time dynamics of the wavefunction globally depends on all Hamiltonian terms. On the other hand, locality is explicit in the Heisenberg picture, for which the time dynamics of local operators only depends on nearby local operators (when the Hamiltonian is local). [1] This motivates us to ask: Is it possible to modify the Schrödinger picture to make locality explicit in the equations of motion?
Gauge theory is another fundamental concept in theoretical physics. Gauge theory is the foundation of the Standard Model of particle physics and is also used to describe exotic phases of condensed matter [2, 3]. An important tool that gauge theory provides is the gauging process, in which one promotes a global symmetry into a local gauge symmetry by coupling the original model to gauge fields. For example, a scalar field theory $\mathcal{L}=\frac{1}{2}(\partial_{\mu}\phi)^{2}$ is invariant under a global $U(1)$ symmetry $\phi(x)\to\phi(x)+\lambda$. By coupling $\phi$ to a gauge field $A_{\mu}$ as in $\mathcal{L}_{\text{gauged}}=\frac{1}{2}(\partial_{\mu}\phi-A_{\mu})^{2}$, the global symmetry is promoted to a local gauge symmetry where $\phi(x)\to\phi(x)+\lambda(x)$ and $A_{\mu}(x)\to A_{\mu}(x)+\partial_{\mu}\lambda(x)$. Gauging more exotic symmetries leads to more exotic physics; e.g. gauging spatial symmetries can lead to gravity and gauging fractal symmetries can result in fracton topological order [4]. In quantum mechanics, expectation values $\braket{\psi}{A}{\psi}$ are invariant under a global unitary transformation acting on the wavefunction |$\ket{\psi}\to U\ket{\psi}$ and operators $A\to UAU^{\dagger}$. Although this transformation is typically viewed as a global invariance rather than a global symmetry, we can still ask: Is it possible to gauge the global unitary invariance in quantum mechanics? And what are the consequences of doing so?
@@ 1 Introduction
The Schrödinger and Heisenberg pictures are related by a time-dependant unitary transformation that acts on the wavefunctions and operators. Our new picture of quantum dynamics is related to the Schrödinger and Heisenberg pictures via a local gauge transformation. We describe these derivations in detail in Sec. 2 and Sec. 3. The new local equations of motion are given in Eq. (19), and the local gauge transformation is given in Eq. (22).
In Sec. 4, we show that the gauge picture local wavefunction associated with a patch is equivalent to the interaction picture wavefunction when the interaction Hamiltonian is the sum over Hamiltonian terms that have some support on that patch. To gain intuition, in Sec. 5 we consider the example of a quantum circuit in the gauge picture. In Sec. 6, we describe the measurement process in the gauge picture. Although local unitary dynamics are explicitly local in the gauge picture, the affect of local measurement is not explicitly local in the gauge picture (similar to the Schrödinger picture). In Sec. 7, we generalize spatial locality in the gauge picture to e.g. locality in local particle number or local energy density.
2 Schrödinger to Heisenberg Picture
@@ 3 Gauge Picture
3.1 Solving the Equations of Motion
@@ 3.2 Generalized Hamiltonians
4 Local Interaction Picture
@@
5 Quantum Circuits
(a) Schrödinger:
|$\ket{\psi^{\text{S}}(t_{\text{f}})}=U\ket{\psi_{0}}$
(b) Heisenberg:
$A^{\text{H}}(t_{\text{f}})=UA(0)\,U^{\dagger}$
(c) local Heisenberg:
$A_{I}^{\text{H}}(t_{\text{f}})=UA_{I}(0)U^{\dagger}$
@@ 8 Conclusion
8.1 New Approximation Technique
Simulating the gauge picture dynamics on a computer can be achieved using a computational time that scales as $O(n_{\text{U}}N^{3}T_{\text{f}}/\delta_{\text{t}})$, where $n_{\text{U}}$ is the number of connections (which typically scales with the system size), $N$ is the Hilbert space dimension, $N^{3}$ is the time complexity for $N\times N$ matrix multiplication in practice [^2] For extremely large $N$, an $N^{\omega}$ time complexity for matrix multiplication is in principle possible for some $32\leq\omega<$ 3 [5, 6]., $T_{\text{f}}$ is the final time, and $\delta_{\text{t}}$ is the time step used for numerical integration. [^3] The time complexity is dominated by the time it takes to multiply a connection $U_{IJ}$ with another $N\times N$ matrix in Eqs. (19) and (20), which must be done $O(n_{\text{U}}T_{\text{f}}/\delta_{\text{t}})$ times. Although this is more expensive than simulating the Schrödinger picture, in another work [7] we show that the gauge picture facilitates a new kind of approximation scheme for improved computational efficiency.
@@ 8.1 New Approximation Technique
Figure 4: (a) A matrix product state in canonical form, where the yellow tensors are isometry tensors. The red tensor can be viewed as a local wavefunction in a truncated Hilbert space. (b) Removing the red tensor from the matrix product state yields the operator that maps the full Hilbert space (bottom) to the truncated Hilbert space (top). For more details, see Appendix B of Ref. [7] (from which this figure was adapted).
@@
The gauge picture allows a more efficient Hilbert space truncation where a different subspace of states is kept for different patches of space. The utility of a position-dependent Hilbert space truncation is exemplified by matrix product states [8] (and isometric tensor networks [9]), which are efficient tensor network representations of wavefunctions with area-law entanglement [8]. Figure 4 (adapted from Ref. [7]) shows how a matrix product state can be viewed as a wavefunction in a position-dependent truncated Hilbert space. Although position-dependent truncated Hilbert spaces are already used by matrix product states and isometric tensor networks, the gauge picture provides a new way to approximately time evolve wavefunctions using position-dependent truncated Hilbert spaces. Compared to tensor network methods, the gauge picture provides significantly simpler quantum dynamics algorithms for models in two or more spatial dimensions and for models with fermions. See Ref. [7] for a more thorough exploration of this approach.
8.2 New Deformation of Quantum Mechanics
@@
The gauge picture agrees with the Schrödinger picture when there is no curvature in the gauge connections, i.e. when Eq. (14) is satisfied. Adding curvature to the initial conditions [i.e. adding violations of Eq. (14)] yields a new kind of deformation of quantum theory, which could yield testable experimental predictions for new physics beyond quantum theory. Considering gauge curvature deformations of quantum theory may be appealing since gauge curvature has been a major theme in the Standard Model of particle physics. Furthermore, this deformation explicitly maintains locality, in contrast to Weinberg’s nonlinear deformations of quantum theory [10] which allows supraluminal communications [11, 12]. However, future research is needed to determine if testable predictions [13] can be made that are consistent with previous experiments.
8.3 Other Future Directions
@@
Other intriguing future directions include: (1) It is not clear how to describe quantum channels or the Lindblad master equation in the gauge picture. The obstruction is that we used a unitary operator ($\widetilde{U}_{I}$) to cancel out distant unitary dynamics in order to achieve locality; but unitary operators can not cancel out the non-unitary dynamics of quantum channels or the Lindbladian. (2) Does the gauge picture have useful applications for understanding exotic dynamical many-body phenomena, such as out-of-time-ordered correlation functions [14, 15], many body localization [16, 17], or the speed of information propagation [18, 19, 20]?